OpenStudy (selena2345):

Determine which equation is an identity or whether it has no solution. a-4a=2a+1-5a

OpenStudy (anonymous):

There is only one equation. Are there supposed to be more?

OpenStudy (selena2345):

a-4a=2a+1-5a @BlackLabel

OpenStudy (anonymous):

This one has no solution

OpenStudy (selena2345):

can u show me how its no solution ? @BlackLabel

OpenStudy (anonymous):

OK well lets solve it by gathering like terms on both sides a-4a=2a+1-5a -3a=2a+1-5a

OpenStudy (anonymous):

Did you follow so far? I just added/subtracted like terms

OpenStudy (selena2345):

yes @BlackLabel

OpenStudy (anonymous):

Ok will do the same thing for the left hand side -3a=2a+1-5a -3a=-3a+1

OpenStudy (selena2345):

Ok @BlackLabel

OpenStudy (anonymous):

Does anything seem weird about -3a=-3a+1 @Selena2345

OpenStudy (selena2345):

No @BlackLabel

OpenStudy (anonymous):

Ok well then lets continue solving lets add 3a to both sides 3a-3a=-3a=3a+1 0=1 CAN 0 EQUAL 1???????

OpenStudy (anonymous):

@Selena2345 ?

OpenStudy (anonymous):

I gotta go soon so any questions?

OpenStudy (selena2345):

yes @BlackLabel

OpenStudy (selena2345):

One More @BlackLabel

OpenStudy (anonymous):

I dont have time now cuz I gotta go out but you can ask someone else if you have anotehr problem

OpenStudy (anonymous):

if ya have a question on what I explained above I can answer quickly

OpenStudy (selena2345):

Really @BlackLabel

OpenStudy (anonymous):

Ok gtg bbl

OpenStudy (selena2345):

9x+3x-10+3(3x+x)@BlackLabel

OpenStudy (selena2345):

Please @BlackLabel